OpenStudy (anonymous):

I want to know is this write Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 5x + 5y = 50 10x + 6y = 84

OpenStudy (anonymous):

5x + 5y = 50 2) 10x + 6y = 84 .. multiply eq 1 by -2 3) -10x-10y=-100 2) 10x + 6y = 84 add -4y=-16 y=4 .. 5x=50-5y=50-20=30 x=30/5=6 one solution, lines intersect at (6,4)

OpenStudy (anonymous):

correct

OpenStudy (angela210793):

tht's right

OpenStudy (akshay_budhkar):

:) right..lol

OpenStudy (anonymous):

ty all im new to this type of math

OpenStudy (anonymous):

Y=4 AND X=6 THE SAME ANS